OpenStudy (_excusemyfrench._):

Fan and Medal..Check my answer? What is the next term in the number pattern? 77,64,53,44,37.... I think its 32

OpenStudy (karatechopper):

Hm Lets look at the difference between each 77-64=13 64-53=11 53-44=9 44-37=7

OpenStudy (karatechopper):

Whats the pattern between the numbers on the right?

OpenStudy (_excusemyfrench._):

the numbers being subtracted go down by 2 each time

OpenStudy (_excusemyfrench._):

so in the last one 7-2=5 and 37-5 32

OpenStudy (karatechopper):

perfect there you go! :)

OpenStudy (_excusemyfrench._):

Thanks!!